OpenStudy (lerenstudy):

How would you find the minimum value of f(x) = 3x^2 + 12x + 16

OpenStudy (across):

What have you tried?

OpenStudy (lerenstudy):

I tried using this formula, because I remembered it, but it came out weird: (c - b^2/4a)

OpenStudy (lerenstudy):

Rather c-(b^2/4a)

OpenStudy (lerenstudy):

I got 4 by doing that, and I wasn't sure if I did that correctly or not

OpenStudy (welshfella):

you can also do it by converting to vertex form . Complete the square.

OpenStudy (across):

That's the right answer. The simplest way is to differentiate: \(f'(x)=0\implies x=-2\implies f(-2)=4\). Your method only gives you the \(y\)-coordinate of the minimum, though: it does not tell you where it happens.

OpenStudy (lerenstudy):

Okay, so then substitute in 4 for the y-value?

OpenStudy (across):

That's right: solve \(4=3x^2+12x+16\).

OpenStudy (lerenstudy):

4 = 3x^2 + 12x + 16 -12=3x^2+12x -4=x^2+3x -4=(x+1.5)^2 -1.75=(x+1.5)^2 Did something go wrong, because I know that you can't square root a negative.

OpenStudy (across):

check third line

OpenStudy (lerenstudy):

Wait, would it be 2 because x^2+4x (because I messed up and wrote 3 instead) +4 and if you use product sum, it would be 2 and 2 because 2*2 =4 and 2+2=4?

OpenStudy (across):

correct

OpenStudy (lerenstudy):

But then you would negate that, correct

OpenStudy (across):

remember that the goal is solving for x.

OpenStudy (lerenstudy):

Right. Would the function would turn into (x+2)(x+2) then?

OpenStudy (across):

correct. you get a repeated root at 2, which means that your min is at (2,4)

OpenStudy (across):

at -2*

OpenStudy (across):

(-2,4)*

OpenStudy (lerenstudy):

Right, okay! Thank you very much!

OpenStudy (mallorysipp234):

simply take the equation and plug it into the y= screen on your calculator and then click 2nd - trace - click minimum, then go to the left and right of your minimum location and when you do that click enter. It will automatically give u it, no work. easier. just gotta remember the method to it.